It’s time to get a new laptop that is $2500. If you save up for it each month it will take one year in an account that earns 5% annual interest. (A)How much would you have to put aside each month to have enough for it? (B)What are the total acquisition costs of saving up for the laptop?

*Please include any formulas used
Answer #1

I = 8% = 8% /12 = 0.6667%

Loan = 2500

t = 2 yrs = 2*12 = 24 months

(A/P,i%,n) = i((1 + i)^n)/((1 + i)^n-1)

Monthly payment = 2500*(A/P,0.6667%,24)

= 2500 * [0.006667*((1 + 0.006667)^24)/((1 + 0.006667)^24-1)]

= 2500 * [0.006667*((1.006667)^24)/((1.006667)^24-1)]

= 2500 * 0.045227

= 113.07

Total acquisition cost = 24 * 113.07 = 2713.68